Circular No. 8311

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004ax AND 2004ay Two apparent supernovae have been reported on unfiltered CCD im-

ages, SN 2004ax by D. Singer and W. Li (LOSS/KAIT; cf. IAUC 8307) and SN 2004ay by T. Boles (cf. IAUC 8290).

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004ax Mar. 21.48 15^{h}24^{m}46.^{s}76 +68^{o}43′56′′.3 17.7 3′′.6 E, 5′′.9 N
2004ay Mar. 23.14 18 28 57.57 +51 38 55.7 18.1 0′′.2 E, 8′′.5 S

Additional approximate KAIT magnitudes for SN 2004ax in NGC 5939: Mar. 6.52 UT, [19.0; 23.53, 17.7. Additional magnitudes by Boles for SN 2004ay in UGC 11255: 2003 Sept. 25, [19.5; Oct. 24, [19.5; 2004 Mar. 24.098, 18.1. SN 2004ay also is not present on Digitized Sky Survey red (1991) and blue images (1990).

SUPERNOVAE 2002lu AND 2002lv T. Schrabback, Institut für Astrophysik und Extraterrestrische For-

schung, Universität Bonn, reports the discovery of two apparent supernovae on archival Hubble Space Telescope ACS i-band images (F775W filter):

SN  2002 UT  α_{2000}  δ_{2000}  i  Offset
2002lu Sept. 23.7 3^{h}33^{m}31.^{s}44 − 9^{o}16′52′′.3 24.0 0′′.4 E
2002lv Sept. 23.7 3 33 33.75 − 9 18 53.0 25.2 0′′.3 E, 0′′.5 N

Additional i magnitudes: 2002lu, 2002 Aug. 11 UT, [26.7; Sept. 26.67, 24.2; 2003 Aug. 11, [26.9. 2002lv, 2002 Aug. 11, [26.2; Sept. 26.67, 25.2; 2003 Aug. 11, [26.4.

SUPERNOVA 2004aw IN NGC 3997 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Cen-

ter for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2004aw (cf. IAUC 8310), obtained by P. Berlind on Mar. 21.29 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a supernova of type I, but the exact classification is uncertain. The spectrum has peaks near 400, 450, 505, 570, and 710 nm; there is an absorption feature near 610 nm. The spectrum shows a strong similarity with both the type-Ia supernova 2001V at day −14 and the peculiar type-Ic supernova 2003jd at early times.

2004 March 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green